write in standard form y=3/4x+6
Answer by mananth(16946)

y = (3/4)x + 6
-6 -6
y-6 = 3x/4
multiply by 4
4(y-6)=3x
4y-24=3x
-3x -3x
-3x+4y-24 =0
-3x+4y=24
multiply by -1
3x-4y=-24
...
m.ananth@hotmail.ca